OpenStudy (anonymous):

Which ratio is proportional to 5/8 ?

OpenStudy (solomonzelman):

\[\frac{5 \times x}{8\times x}\]

OpenStudy (anonymous):

7/10 15/18 25/45 45/72

OpenStudy (anonymous):

10/16 20/32 40/64 80/128

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

lee those number r wrong the ones u put

OpenStudy (anonymous):

Those are the answers